OpenStudy (anonymous):

Write and then solve for y = f(x) the differential equation for the statement: "The rate of change of y with respect to x is inversely proportional to y^4."

OpenStudy (michele_laino):

hint: two quantities, say A and B are inversely proportional if the subsequent condition olds: A*B = k where k is a constant

OpenStudy (michele_laino):

holds*

OpenStudy (anonymous):

dy/dx = ky^4 ?

OpenStudy (michele_laino):

no, since that formula expresses the direct proportionality

OpenStudy (michele_laino):

hint: we can write this: A= dy/dx, and B=y^4

OpenStudy (anonymous):

dy/dx = k/y^4

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

okay and now int(dy y^4) = int(dx k) (y^5)/5 = (x^2)/2 + C

OpenStudy (anonymous):

\[y = \sqrt[5]{\frac{ 5x^2 }{ 2 } + C}\]

OpenStudy (michele_laino):

I got this: \[\Large \begin{gathered} \int {{y^4}dy} = \int {kdx} \hfill \\ \hfill \\ \frac{{{y^5}}}{5} = kx + C \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

doesn't the right side turn into \[\frac{ kx^2 }{ 2 } + C\]

OpenStudy (anonymous):

no wait i see what i did wrong

OpenStudy (anonymous):

okay so then \[y = \sqrt[5]{5kx + C}\]

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

sorry! I had to specify this: \[\Large C \in \mathbb{R}\] since you are searching for a real funtion

OpenStudy (anonymous):

thank you so much :D

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

quick question what does the \[\epsilon \] mean

OpenStudy (michele_laino):

in general, in mathematical analysis, the mathematicians use \epsilon in order to indicate a small and positive quantity, which, in some case goes to zero

OpenStudy (anonymous):

okay , double thank you :D

OpenStudy (michele_laino):

:)